OpenStudy (anonymous):

solve: 6^6x=124 Round to the nearest ten-thousandth

OpenStudy (campbell_st):

do you know about logs...?

OpenStudy (anonymous):

Kinda sorta ! I got 0.8034 , am i right??

OpenStudy (anonymous):

I just am not sure how to round to the nearest ten thousandth

OpenStudy (campbell_st):

nope here is my solution take the base 6 log of both sides using the log law \[\log _{a}a^b = b\] then \[\log _{6}6^x = \log _{6}(124)\] or \[6x = \log _{6}(124)\] next use change of base to base e logs on the right hand side \[6x = \frac{\ln(124)}{\ln(6)}\] 6x = 2.69025 divide by 6 gives x = 448375 you'll need to round the answer

OpenStudy (anonymous):

I diiiiiid it (: thank you SOooooo much !

OpenStudy (campbell_st):

oops I forgot a decimal point x = 0.448375